# HyGEN: Regularizing Negative Hyperedge Generation for Accurate Hyperedge Prediction

### v1

@@ Abstract.
Hyperdge prediction, Adversarial learning, Negative sampling
@@ 1. Introduction
A hypergraph, a generalized data structure, is capable of modeling a high-order (i.e., group-wise) relation among an arbitrary number of real-world objects as a hyperedge. Due to its powerful expressiveness, hypergraph-based network learning (et al., 2019a, 2022b) has been widely studied and shown to outperform graph-based methods in various downstream tasks, including node classification (et al., 2020b, 2023a), node ranking (et al., 2021c, b), and link prediction (et al., 2020e; Vaida and Purcell, 2019).
@@
Hyperedge prediction (i.e., link prediction on hypergraphs) is a fundamental task in many real-world applications, such as recommender systems (et al., 2021c) and social network analysis (et al., 2021b, 2023a); it predicts future high-order relations (i.e., hyperedges) based on an observed hypergraph structure. A general approach to hyperedge prediction is two-fold (et al., 2022a, 2023b) (see Figure 1: (1) (hypergraph encoding) the embeddings of nodes are produced by hypergraph neural networks (et al., 2019a, 2020b) and (2) (hyperedge candidate scoring) the embeddings of nodes in each hyperedge candidate are aggregated and fed into a predictor to determine whether the candidate is real.
@@
However, sampling ‘good’ negative hyperedges is challenging in the context of hyperedge prediction since there exist too many possible negative hyperedges (i.e., $2^{|V|}-|E|$). Although existing methods (et al., 2019b, 2020e), enhanced by NS, have achieved breakthroughs in many fields, they focus primarily on hypergraph encoding while employing simple heuristic-based NS methods (et al., 2020c). Thus, negative sampling for hyperedge prediction is still rarely explored.
Although one recent work (et al., 2022a) proposed an adversarial-training-based hyperedge prediction method (AHP) that leverages model-generated negative hyperedges useful for model training; however, it has overlooked the following two important challenges:
(C1) Lack of guidance for generating negatives. Only a random noise is used as the input of the hyperedge generator in AHP. Thus, it may fail to effectively reflect the characteristics of positive hyperedges into generated hyperedges, especially in the initial stage of the training, which could lead to inefficient and unstable training.
(C2) Possibility of producing false negatives. AHP generates negative hyperedges by using a generative adversarial network (GAN), which aims to generate negative hyperedges similar as much as possible to positive hyperedges (i.e., copying the exact distribution of the original hyperedges). Without any regularization, however, it might be possible to generate hyperedges too similar to positives, which would be potential positive hyperedges.
@@ 2. Related Works
Hyperedge prediction. There have been a number of works to study hyperedge prediction. Expansion (et al., 2020d) models a hypergraph as multiple n-projected graphs and predicts future hyperedges based on the projected graphs. HyperSAGNN (et al., 2019b) uses a self-attention-based GNN model to learn hyperedges of variable sizes. NHP (et al., 2020e) employs hyperedge-aware GNN models to learn node embeddings in hypergraphs, using the max-min pooling to aggregate the embeddings of nodes within each hyperedge candidate for prediction. AHP (et al., 2022a), the state-of-the-art hyperedge prediction method, employs adversarial training to generate negative hyperedges for model training and uses max-min pooling for node aggregation.
@@ 3. Proposed Method: HyGEN
Notations. A hypergraph is defined as $H=(V,E)$, where $V=\{v_{1},v_{2},\dots,v_{|V|}\}$ and $E=\{e_{1},e_{2},\dots,e_{|E|}\}$. A hypergraph can generally be represented by an incidence matrix $\mathbf{H}\in\{0,1\}^{|V|\times|E|}$, where each element $h_{ij}=1$ if $v_{i}\in e_{j}$, and $h_{ij}=0$ otherwise. The node and hyperedge features are represented by the matrices $\mathbf{P}\in\mathbb{R}^{|V|\times d}$, $\mathbf{Q}\in\mathbb{R}^{|E|\times d}$, where each row $p_{i}$ and $q_{i}$ represents the $d$-dimensional feature of a node and a hyperedge, respectively.
@@
(3) Model training. We train the model parameters of HyGEN in an adversarial way (et al., 2017a): given a batch $B$ of positive hyperedges, (1) generate $|B|$ negative hyperedges using the generator $G$ ($enc(\cdot)$ and $dec(\cdot)$), (2) classify the positive and negative hyperedges using the discriminator $D$ ($agg(\cdot)$ and $pred(\cdot)$), and (3) update the model parameters of HyGEN based on their losses. Specifically, as $D$ aims to compute the probabilities of positive hyperedges higher than those of negative hyperedges, the loss function for $D$ is defined as:
@@
where $k$ and $p$ are hyperparameters that control the converge point and the curvature of the function, respectively (See Figure 3). $sim(\cdot)$ is the cosine similarity used as a similarity function. Finally, the regularization loss is integrated into the total loss with a hyperparameter $\beta$ to control the weight of the regularization loss:
@@
Figure 3. Regularization loss according to hyperparameters.
@@ 4. Experimental Validation
We evaluate HyGEN by answering the following questions.
@@ 4.1. Experimental Setups
Datasets and competitors111 We have released all the code and datasets at: https://github.com/ssong915/HyGEN. We use six widely used real-world hypergraphs: (1) three co-citation datasets (Citeseer, Cora, and Pubmed), (2) two authorship datasets (Cora-A and DBLP-A), and (3) one collaboration dataset (DBLP). We select four state-of-the-art hyperedge prediction methods as our competitors (Expansion (et al., 2020d), NHP (et al., 2020e), HyperSAGNN (et al., 2019b), and AHP (et al., 2022a)).
Evaluation protocol. We evaluate HyGEN by using the protocol exactly same as that used in (et al., 2022a). For each dataset, we use five data splits, where positive hyperedges are randomly divided into training (60%), validation (20%), and test (20%) sets. We use three validation and test sets constructed with negative hyperedges sampled by SNS, MNS, and CNS, explained in Section 2. As metrics, we use AUROC (area under the ROC curve) and AP (average precision). We (1) measure AUROC and AP on each test set when the averaged AUROC over the validation sets is maximized, and (2) report the averaged AUROC and AP over five runs on each test set.
AUROC (%) AP (%) AUROC (%) AP (%) AUROC (%) AP (%)
SNS MNS CNS AVG SNS MNS CNS AVG SNS MNS CNS AVG SNS MNS CNS AVG SNS MNS CNS AVG SNS MNS CNS AVG
Expansion Citeseer 66.3 78.1 33.1 59.2 76.5 81.7 49.8 69.3 Pubmed 52.0 73.0 24.1 49.7 67.5 75.5 44.0 62.3 Cora 47.0 70.7 25.6 47.8 63.7 76.4 45.4 61.8
NHP 99.1 70.1 51.0 73.4 99.0 73.1 52.0 74.7 97.3 69.4 52.4 73.0 97.3 65.6 51.3 71.4 94.3 64.1 47.2 68.5 94.9 67.8 50.9 71.2
HyperSAGNN 54.0 41.0 47.3 47.4 62.7 45.5 49.7 52.6 52.5 68.6 54.6 58.6 53.4 68.0 52.9 58.1 61.7 52.7 49.4 54.6 68.7 57.4 50.8 59.0
AHP 94.3 88.1 65.1 82.5 95.2 87.0 66.0 82.7 91.7 84.0 53.3 76.3 91.8 83.4 52.6 75.9 96.4 86.0 57.2 79.9 96.1 83.7 55.2 78.3
HyGEN 98.4 92.6 67.6 86.2 98.5 91.2 69.4 86.4 92.1 87.1 51.6 77.0 93.2 89.0 55.1 79.1 99.1 90.7 58.4 82.7 99.0 89.6 57.0 81.9
Expansion Cora-A 69.0 84.2 43.4 65.5 69.0 87.6 57.7 71.4 DBLP 64.5 80.1 36.6 60.4 75.1 85.6 51.8 70.8 DBLP-A 63.4 82.6 35.0 60.3 73.0 85.2 51.2 69.8
NHP 90.9 67.2 55.0 71.0 92.5 72.0 58.5 74.3 66.3 54.0 50.3 56.9 60.8 52.3 50.1 54.4 96.6 62.3 55.5 71.5 96.5 60.4 53.4 70.1
HyperSAGNN 38.6 59.1 54.2 50.6 53.2 64.3 54.5 57.3 44.8 57.4 57.2 53.1 56.2 60.2 58.6 58.3 54.8 79.1 56.3 63.4 68.6 80.5 55.2 68.1
AHP 95.8 92.4 78.2 88.8 95.7 89.8 79.6 88.4 94.6 82.0 56.8 77.8 94.7 81.5 56.1 77.4 91.6 92.6 66.8 83.7 92.8 92.8 70.7 85.4
HyGEN 97.7 94.7 80.3 90.9 97.7 91.4 82.9 90.7 95.8 83.9 59.1 79.6 95.8 82.6 57.3 78.6 91.3 92.7 67.1 83.7 92.7 92.9 71.5 85.7
@@
4.2. Experimental Results
EQ1. Hyperedge prediction accuracy. We compare HyGEN with four competing methods in the hyperedge prediction task. Table 1 shows that HyGEN consistently outperforms all competing methods in terms of both the averaged AP and AUROC over three test sets across all datasets. We note that these improvements of HyGEN over AHP (the best competitor) are remarkable, given that AHP has already improved other existing methods significantly in those datasets. Via the t-tests with a 95% confidence level, we verify that the improvements of HyGEN are statistically significant (i.e., the p-values $\leq 0.05$). As a result, these results demonstrate that HyGEN can generate informative negative hyperedges by effectively addressing the two challenges of negative sampling, thereby enhancing the accuracy of a hyperedge prediction task.
Figure 4. The hyperparameter sensitivity of HyGEN.
EQ2. Ablation study. We verify the effectiveness of our proposed strategies individually by ablating one of them: (i) positive-guided negative hyperedge generator and (ii) regularization term. Table  2 shows that the original version of HyGEN always achieves the highest accuracy across all datasets, which indicates that ablating one of our proposed strategies could lead to accuracy degradation. These results verify the effectiveness of our proposed strategies for addressing the two challenges: (C1) lack of guidance for generating negatives and (C2) possibility of false negatives.
EQ3. Sensitivity analysis. We evaluate the impact of hyperparameters $k$ and $p$ on the accuracy of HyGEN. We measure the model accuracy of HyGEN with varying $k$ from 0 to 1.0 in step of 0.1 and $p$ from 1 to 5 in step of 1. As shown in Figure 4, where the $x$-axis represents the converge point hyperparameter $k$, the $y$-axis represents the curvature hyperparameter $p$, and the $z$-axis represents the averaged AUROC, HyGEN with $k\geq 0.4$ consistently achieves higher accuracy than HyGEN with $k<0.4$ regardless of $p$ (i.e., the red wide area on the surface in Figure 4). Based on these results, we believe that the accuracy of HyGEN is insensitive to the regularization hyperparameters $k$ and $p$.
@@ 5. Conclusion
In this paper, we identified two key challenges of negative hyperedge sampling in the hyperedge prediction task: (C1) lack of guidance for generating negatives and (C2) possibility of producing false negatives. To address both challenges, we propose a novel hyperedge prediction method, HyGEN that employs (1) a positive-guided negative hyperedge generator leveraging positive hyperedges as a guidance to generate informative negative hyperedges for (C1) and (2) a regularization term to prevent the generated hyperedges from being false negatives (C2). Comprehensive experiments on six real-world datasets verified the superiority of HyGEN over four state-of-the-art hyperedge prediction methods.

### v2

@@ Abstract.
Hyperdge prediction, Adversarial learning, Negative hyperedge generation, Regularization
@@ 1. Introduction
In real-world networks, high-order relations (i.e., group-wise relations) are prevalent (et al., 2025, 2020b, 2019a), such as (i) a research paper co-authored by a group of researchers and (ii) a chemical reaction co-induced by a group of proteins. A hypergraph, a generalized data structure, is capable of modeling such group-wise relations as a hyperedge without any information loss. Due to its powerful expressiveness, hypergraph-based network learning (et al., 2019a, 2022b) has been widely studied and shown to outperform graph-based methods in various downstream tasks, including node classification (et al., 2020b, 2023), node ranking (et al., 2021c, b), and link prediction (et al., 2020e; Vaida and Purcell, 2019).
@@
Hyperedge prediction (i.e., link prediction on hypergraphs) is a fundamental task in many real-world applications, such as recommender systems (et al., 2021c) and social network analysis (et al., 2021b, 2023); it predicts future high-order relations (i.e., hyperedges) based on an observed hypergraph structure. A general approach to hyperedge prediction is two-fold (et al., 2022a, 2025) (see Figure 1: (1) (hypergraph encoding) the embeddings of nodes are produced by hypergraph neural networks (et al., 2019a, 2020b) and (2) (hyperedge candidate scoring) the embeddings of nodes in each hyperedge candidate are aggregated and fed into a predictor to determine whether the candidate is real.
@@
However, sampling ‘good’ negative hyperedges is challenging in the context of hyperedge prediction since there exist too many possible negative hyperedges (i.e., $2^{|V|}-|E|$). Although existing hyperedge prediction methods (et al., 2019b, 2020e, 2020d), enhanced by negative sampling methods, have achieved breakthroughs in many fields, they focus primarily on hypergraph encoding while employing simple heuristic-based negative sampling methods (et al., 2020c). Thus, negative sampling for hyperedge prediction is still rarely explored.
Although one recent work (et al., 2022a) proposed an adversarial-training-based hyperedge prediction method (AHP) that leverages model-generated negative hyperedges useful for model training, it has overlooked the following two important challenges:
(C1) Lack of guidance for generating negatives. Only a random noise signal is used as the input of the hyperedge generator in the previous method. Thus, we posit that it may fail to effectively reflect the characteristics of positive hyperedges into generated hyperedges, especially in the initial stage of the training, which could lead to inefficient and unstable training.
(C2) Possibility of producing false negatives. AHP generates negative hyperedges by using a generative adversarial network (GAN), which aims to generate negative hyperedges similar as much as possible to positive hyperedges (i.e., copying the exact distribution of the original hyperedges). Without any regularization for hyperedge generation, however, it might be possible to generate hyperedges too similar to positives, which would be potential positive hyperedges.
@@
For reproducibility, we have released the code of HyGEN and datasets at: https://github.com/ssong915/HyGEN.
@@ 2. Related Works
Hyperedge prediction. There have been a number of works to study hyperedge prediction. they solve the hyperedge prediction problem as a classification task (et al., 2020d, 2019b, e, 2022a). Expansion (et al., 2020d) models a hypergraph as multiple n-projected graphs and applies a logistic regression model to the projected graphs for predicting future hyperedges HyperSAGNN (et al., 2019b) uses a self-attention-based graph neural networks (GNN) model to learn hyperedges of variable sizes. NHP (et al., 2020e) employs hyperedge-aware GNN models to learn node embeddings in hypergraphs, using the max-min pooling to aggregate the embeddings of nodes within each hyperedge candidate for prediction. AHP (et al., 2022a), the state-of-the-art hyperedge prediction method, employs adversarial training to generate negative hyperedges for model training and uses max-min pooling for node aggregation.
@@
Table 1. Notations and their descriptions
Notation Description
$H$ a hypergraph that consists of nodes and hyperedges
$V,E$ the set of nodes, the set of hyperedges
$\mathbf{H}$ the incidence matrix of $H$
$\mathbf{X}$ the input node features
$\mathbf{P},\mathbf{Q}$ the node and hyperedge representations
$f(\cdot)$ a hypergraph encoder
$agg(\cdot)$ a node aggregator for hyperedge candidates
$pred(\cdot)$ a hyperedge predictor
$G(\cdot),D(\cdot)$ the negative hyperedge generator and discriminator
$enc(\cdot),dec(\cdot)$ the encoder and decoder of the generator $G(\cdot)$
$\mathcal{L}(\cdot)$ loss function
$\mathbf{W},b$ the learnable weight and bias matrices
@@ 3. Proposed Method: HyGEN
In this section, we present a novel hyperedge prediction method, named as HyGEN, for accurate hyperedge prediction.
3.1. Problem Definition
Notations. The notations used in this paper are described in Table 1. A hypergraph is defined as $H=(V,E)$, where $V=\{v_{1},v_{2},\dots,v_{|V|}\}$ and $E=\{e_{1},e_{2},\dots,e_{|E|}\}$. A hypergraph can generally be represented by an incidence matrix $\mathbf{H}\in\{0,1\}^{|V|\times|E|}$, where each element $h_{ij}=1$ if $v_{i}\in e_{j}$, and $h_{ij}=0$ otherwise. The node and hyperedge features are represented by the matrices $\mathbf{P}\in\mathbb{R}^{|V|\times d}$, $\mathbf{Q}\in\mathbb{R}^{|E|\times d}$, where each row $p_{i}$ and $q_{i}$ represents the $d$-dimensional feature of a node and a hyperedge, respectively.
@@ 3.2. Methodology
3.3. Model Training
We train the model parameters of HyGEN in an adversarial way (et al., 2017a): given a batch $B$ of positive hyperedges, (1) generate $|B|$ negative hyperedges using the generator $G$ ($enc(\cdot)$ and $dec(\cdot)$), (2) classify the positive and negative hyperedges using the discriminator $D$ ($agg(\cdot)$ and $pred(\cdot)$), and (3) update the model parameters of HyGEN based on their losses. Specifically, as $D$ aims to compute the probabilities of positive hyperedges higher than those of negative hyperedges, the loss function for $D$ is defined as:
@@ 3.3. Model Training
Figure 3. Regularization loss according to hyperparameters.
AUROC AP AUROC (%) AP (%)
SNS MNS CNS AVG SNS MNS CNS AVG SNS MNS CNS AVG SNS MNS CNS AVG
Expansion Citeseer 66.3 78.1 33.1 59.2 76.5 81.7 49.8 69.3 Pubmed 52.0 73.0 24.1 49.7 67.5 75.5 44.0 62.3
NHP 99.1 70.1 51.0 73.4 99.0 73.1 52.0 74.7 97.3 69.4 52.4 73.0 97.3 65.6 51.3 71.4
HyperSAGNN 54.0 41.0 47.3 47.4 62.7 45.5 49.7 52.6 52.5 68.6 54.6 58.6 53.4 68.0 52.9 58.1
AHP 94.3 88.1 65.1 82.5 95.2 87.0 66.0 82.7 91.7 84.0 53.3 76.3 91.8 83.4 52.6 75.9
HyGEN 98.4 92.6 67.6 86.2 98.5 91.2 69.4 86.4 92.1 87.1 51.6 77.0 93.2 89.0 55.1 79.1
Expansion Cora 47.0 70.7 25.6 47.8 63.7 76.4 45.4 61.8 Cora-A 69.0 84.2 43.4 65.5 69.0 87.6 57.7 71.4
NHP 94.3 64.1 47.2 68.5 94.9 67.8 50.9 71.2 90.9 67.2 55.0 71.0 92.5 72.0 58.5 74.3
HyperSAGNN 61.7 52.7 49.4 54.6 68.7 57.4 50.8 59.0 38.6 59.1 54.2 50.6 53.2 64.3 54.5 57.3
AHP 96.4 86.0 57.2 79.9 96.1 83.7 55.2 78.3 95.8 92.4 78.2 88.8 95.7 89.8 79.6 88.4
HyGEN 99.1 90.7 58.4 82.7 99.0 89.6 57.0 81.9 97.7 94.7 80.3 90.9 97.7 91.4 82.9 90.7
Expansion DBLP 64.5 80.1 36.6 60.4 75.1 85.6 51.8 70.8 DBLP-A 63.4 82.6 35.0 60.3 73.0 85.2 51.2 69.8
NHP 66.3 54.0 50.3 56.9 60.8 52.3 50.1 54.4 96.6 62.3 55.5 71.5 96.5 60.4 53.4 70.1
HyperSAGNN 44.8 57.4 57.2 53.1 56.2 60.2 58.6 58.3 54.8 79.1 56.3 63.4 68.6 80.5 55.2 68.1
AHP 94.6 82.0 56.8 77.8 94.7 81.5 56.1 77.4 91.6 92.6 66.8 83.7 92.8 92.8 70.7 85.4
HyGEN 95.8 83.9 59.1 79.6 95.8 82.6 57.3 78.6 91.3 92.7 67.1 83.7 92.7 92.9 71.5 85.7
@@
where $k$ and $p$ are hyperparameters that control the converge point and the curvature of the function, respectively. Figure 3 shows that regularization loss according to the hyperparameters $k$ and $p$. $sim(\cdot)$ is the cosine similarity used as a similarity function. Finally, the regularization loss is integrated into the total loss with a hyperparameter $\beta$ to control the weight of the regularization loss:
@@
Thus, all the model parameters of HyGEN are trained jointly based on the adversarial loss and the regularization loss.
@@ 4. Experimental Validation
In this section, we comprehensively evaluate HyGEN by answering the following evaluation questions (EQs):
@@ 4.1. Experimental Setups
Datasets and competitors We use six widely used real-world hypergraphs: (1) three co-citation datasets (Citeseer, Cora, and Pubmed), (2) two authorship datasets (Cora-A and DBLP-A), and (3) one collaboration dataset (DBLP). In the co-citation datasets, each node represents a paper and each hyperedge represents a group of papers co-cited by a paper; in the authorship dataset, each node represents a paper and each hyperedge represents a group of papers written by an author; in the collaboration dataset, each node represents a researcher and each hyperedge represents a group of researchers who wrote the same paper. For all the datasets, we use the bag-of-word features from the abstract of each paper as in (et al., 2022a, 2025). We select four state-of-the-art hyperedge prediction methods as our competitors in the experiments (Expansion (et al., 2020d), NHP (et al., 2020e), HyperSAGNN (et al., 2019b), and AHP (et al., 2022a)).
Evaluation protocol. We evaluate HyGEN by using the protocol exactly same as that used in (et al., 2022a). For each dataset, we use five data splits  [^1] All datasets and their splits used in this paper are available at: https://github.com/ssong915/HyGEN., where positive hyperedges are randomly divided into training (60%), validation (20%), and test (20%) sets. We use three validation and test sets constructed with negative hyperedges sampled by SNS, MNS, and CNS, explained in Section 2. As metrics, we use AUROC (area under the ROC curve) and AP (average precision). We (1) measure AUROC and AP on each test set when the averaged AUROC over the validation sets is maximized, and (2) report the averaged AUROC and AP over five runs on each test set. For all competing methods, we use the results reported in (et al., 2022a) since we follow the exactly same protocol with the same data splits.
@@
Figure 4. Sensitivity analysis: HyGEN achieves high accuracy across a wide range of values for the hyperparameters $k$ and $p$.
4.2. EQ1. Hyperedge Prediction Accuracy
We first compare HyGEN with four competing methods in the hyperedge prediction task. Table 2 shows that HyGEN consistently outperforms all competing methods in terms of both the averaged AP and AUROC over three test sets across all datasets. We note that these improvements of HyGEN over AHP (the best competitor) are remarkable, given that AHP (et al., 2022a), the state-of-the-art method. has already improved other existing methods significantly in those datasets. Via the t-tests with a 95% confidence level, we verify that the improvements of HyGEN are statistically significant (i.e., the p-values $\leq 0.05$). As a result, these results demonstrate that HyGEN can generate informative negative hyperedges by effectively addressing the two challenges of negative sampling, thereby enhancing the accuracy of a hyperedge prediction task.
Interestingly, NHP (et al., 2020e) achieves the highest accuracies in the SNS test setting of the Citeseer, Pubmed, and DBLP-A, However, NHP shows very low accuracies on the CNS test set (i.e., the most difficult test set) of all datasets, which is similar to or even worse than the accuracy of the random prediction ($\approx 0.5$), These accuracy gaps between the CNS and SNS test sets imply that NHP may be overfitting to the easy negative examples, thus which limits their ability to be generalized to other datasets. On the other hand, HyGEN consistently achieves very high accuracies across all test test settings, for example 98.4%, 92.1%, and 91.3% in the SNS setting of Citeseer, Pubmed, and DBLP-A, respectively.
4.3. EQ2. Ablation Study
We verify the effectiveness of our proposed strategies of HyGEN individually by ablating one of them: (i) positive-guided negative hyperedge generator and (ii) regularization term. Table  3 shows that the original version of HyGEN always achieves the highest accuracy across all datasets, which indicates that each of the proposed strategies is always beneficial to improving the accuracy of HyGEN. These results verify that our proposed strategies are able to address the two challenges successfully: (C1) lack of guidance for generating negatives and (C2) possibility of false negatives.
Furthermore, HyGEN w/o $L_{reg}$ shows the worst results in all cases, indicating that ablating the regularization term can lead to significant accuracy degradation. These results verify that HyGEN can generate negative hyperedges that are sufficiently distinct from positive hyperedges, making them useful for model training.
4.4. EQ3. Sensitivity Analysis
In this experiment, we evaluate the impacts of regularization hyperparameters $k$ and $p$ on the accuracy of HyGEN. We measure the model accuracy of HyGEN with varying $k$ from 0 to 1.0 in step of 0.1 and $p$ from 1 to 5 in step of 1. Figure 4 shows the results, where the $x$-axis represents the converge point hyperparameter $k$, the $y$-axis represents the curvature hyperparameter $p$, and the $z$-axis represents the averaged AUROC. HyGEN with $k\geq 0.4$ consistently achieves higher accuracy than HyGEN with $k<0.4$ regardless of $p$ (i.e., the wide red area on the surface in Figure 4). Based on these results, we believe that the accuracy of HyGEN is insensitive to the regularization hyperparameters $k$ and $p$ provided that $k\geq 0.4$.
@@ 5. Conclusion
In this paper, we identify two key challenges of negative hyperedge sampling in the hyperedge prediction task: (C1) lack of guidance for generating negatives and (C2) possibility of producing false negatives. To address both challenges, we propose a novel hyperedge prediction method, HyGEN that employs (1) a positive-guided negative hyperedge generator leveraging positive hyperedges as a guidance to generate informative negative hyperedges for (C1) and (2) a regularization term to prevent the generated hyperedges from being false negatives (C2). Comprehensive experiments on six real-world datasets verified the superiority of HyGEN over four state-of-the-art hyperedge prediction methods.